€2.7, 3.1-3.2 Problem 3.07 MacBook
く2.7, 3.1-3.2 Problem 3.07 Part C Submit Part D
Answer #1

(a) Vz component of velocity will be zero since there is no acceleration in z direction.

(b) Acceleration is in direction 34 degrees above the x axis.

So y component of acceleration will be = a * Sin(34) = 4*(0.56) = 2.23 m/sec2

so Vy = Uy + Ay*t

Vy = 0 + 2.23*22 = 49.21 m

(c) dx position of the particle : from the equation of motion ,

dx = Ux*t + (1/2)*a*t^2

Ux = 0

t = 22 sec

a = Ax = A*cos(34) = 3.3 m/sec2

dx = (1/2)*3.3*22*22 = 798.6 m

(d) Similarly dy position will be

dy = Uy*t + (1/2)*a*t^2

dy = (1/2)*2.23*22*22 = 539.66 m